A hydrocarbon,W with a mass 10.50 g has 17.35% hydrogen. W

undergoes complete combustion.


Calculate the molecular formula of W with the molar mass

58.12gmol -1 .


1
Expert's answer
2021-09-06T02:38:04-0400

This hydrocarbon contains 100 - 17.35 = 82.65% of carbon.

M (H) = 1 g/mol

M (C) = 12 g/mol

H = (0.1735 x 58.12) / 1 = 10

C = (0.8265 x 58.12) / 12 = 4

The molecular formula of W is C4H10.